For the data provided in Question 5, what is the % error for the student’s data?
Answer #1

ERROR= 0.2541-0.2542 =-0.0001

RELAATIVE ERROR = ERROR/ACTUAL VALUE = 0.0001/0.2542 = 0.000393

% ERROR = RELATIVE ERROR *100 = 0.000393*100 =0.0393%

FOR RUN-2

ERROR= 0.2668-0.2542 =0.0126

RELAATIVE ERROR = ERROR/ACTUAL VALUE = 0.0126/0.2542 = 0.0495

% ERROR = RELATIVE ERROR *100 = 0.0495*100 =4.95%

FOR RUN-3

ERROR= 0.2549-0.2542 =0.0007

RELAATIVE ERROR = ERROR/ACTUAL VALUE = 0.0007/0.2542 = 0.0027

% ERROR = RELATIVE ERROR *100 = 0.0027*100 =0.27%

FOR RUN-4

ERROR= 0.2554-0.2542 =0.0012

RELAATIVE ERROR = ERROR/ACTUAL VALUE = 0.0012/0.2542 = 0.0047

% ERROR = RELATIVE ERROR *100 = 0.0047*100 =0.47%